# BEADs: Bias Evaluation Across Domains

### v2

@@ 3.1 Text Classification for Bias, Toxicity, and Sentiment
We fine-tuned smaller BERT-like models for the task. We also fine-tuned autoregressive LLMs including Llama-2-7B-chat [73] and Mistral-7B-Instruct-v0.2 [37] using QLoRA [17] (Appendix G.3), and evaluated them in standard and few-shot setups. Unlike BERT-like models that output probability scores, these LLMs produce categorical outputs, so we analyzed them for precision, recall, and accuracy as in [35]. We also report performance on out-of-distribution datasets (Appendix G.2) for bias, toxicity, and sentiment in Table 22.
Results: As shown in Table 5, fine-tuning smaller BERT-like models can outperform LLMs on specific bias classification tasks, demonstrating their efficiency and practicality for real-world applications. Fine-tuned BERT-like models achieved approximately 85% accuracy in text classification, surpassing autoregressive LLMs in fine-tuning as well as few-shot setups.
@@ 3.2 Token Classification Task
For the token classification task, we fine-tuned models including BiLSTM-CRF, BERT-like models, and LLMs like Llama2-7B-chat and Mistral2-7B-instruct. These models were evaluated on their ability to classify words as ‘Bias’ or ‘O’ (outside). We present this portion of data as CoNLL Bias Dataset following CONLL-2003 format [65] standards for NER, where ‘B-BIAS’ and ‘I-BIAS’ tags are consolidated into ‘Bias’, and ‘O’ into ‘Non-Bias’ for evaluation.
@@
Results: As shown in Table 6, the smaller fine-tuned BERT-like models achieved the highest scores across all metrics, in classifying both ‘Non-Bias’ and ‘Bias’ entities. Few-shot models like Llama2-7B and Mistral-7B delivered moderate performance, where as fine-tuning LLMs tend to show increased performance for token classification task. We analyze aspects of bias in our work is meant, with details in Appendix F.2 and H.2.
@@ 3.3 Bias Quantification on Demographics
Task Setting This task assesses language models’ predictions of demographic identifiers (gender, race, religion) in sentences. Identified mentions in original texts are templated with placeholders like RACE and RELIGION. Format Sentences include placeholders for demographic indicators, structured as <text with identity mention> (e.g., “The {identity mention} individual performs well in managerial jobs"), as in our Bias Quantification Demographics Dataset.
We evaluated both smaller BERT-like models and autoregressive LLMs by testing their ability to fill in placeholders. Their performance was assessed based on Neutrality Rate (alignment with neutral labels) and Bias Rate (alignment with biased labels), details in Appendix G.6.3. The comparison with ground truth bias labels is presented in Table 7 for Neutrality Rates, and for Bias Rates in Table 23.
@@
Table 7: Neutrality Rates for Identity Mentions Across Different Models. The table only displays neutrality rates as the bias rate is the complement to 100%. Higher neutrality rates ($\uparrow$) are desired, with lower implicit bias rates. Models exacerbating bias are highlighted in red.
Identity Mention Original DistilBERT BERT RoBERTa Mistral-7B-v0.2 ethnicity black 33.3 53.6 46.2 56.8 50.0 ethnicity white 42.9 22.2 42.9 45.5 50.0 religion christian 72.4 66.7 62.3 48.9 73.7 religion jewish 50.0 36.7 36.8 40.4 50.0 religion muslim 60.0 43.4 42.2 44.3 50.0 gender female 42.3 40.3 50.8 42.9 43.6 gender male 54.0 47.3 44.7 46.7 57.0
@@
Task Setting: This task evaluates how language models respond to neutral sentences with respect to various demographic groups. We curated 520 neutral sentences from our dataset and created variations for gender (male, female), race (Asian, Black, Hispanic, White), and religion (Christian, Muslim, Hindu, Buddhist). The goal is to determine if the model considers these sentences biased or neutral, when assigned with a certain demographic. Format: Entries are structured as <prompt><text><demographic group variation>, e.g., “Asians are good at math….." or “Europeans are good at maths…..” as in our Demographic Variation Dataset.
LLMs like Mistral-7B-Instruct-v0.2 and Llama2-7B-chat are used during inference to identify and respond to sentences with demographic variation in the prompts. The models responses to these prompts reveal their tendencies to misclassify neutral statements as biased, indicating potential biases toward specific demographics.
@@
Results: The analysis presented in Figure 3 indicates that both Llama2-7B-chat and Mistral-7B-Instruct-v0.2 generally classify neutral sentences as “Neutral” across most demographic categories, which suggests high accuracy. However, there are notable exceptions in categories like “Black”, “Hispanic", “Muslim", and “Christian", where a higher proportion of “Biased” responses might reflect the models heightened sensitivity to or inherent biases in these demographics. Discussion on the result of Stereotype Alignment task on Stereotype Prompts Dataset is in Appendix H.2.
@@ 3.4 Language Generation Task
Task Setting: This task generates neutral or benign versions of biased texts, preserving context while removing bias. The dataset portion has 8,300 entries with various demographics and bias aspects. Format: Each sentence pair is presented as <original sentence>, <benign variation>. For example, Original Sentence: “The young woman was too emotional to be taken seriously at the meeting", Benign Sentence: “The young woman’s passionate response was not fully considered at the meeting". The dataset is our Language Generation Dataset. We also format this dataset portion into Alpaca [72] format for instruction-fine-tuning.
Table 8: Comparison of Bias, Toxicity, Knowledge Retention (KR), Faithfulness (Faith.), and Answer Relevancy (Rel.) across different model variations (Llama2-7B-chat and Mistral-7B-instruct-v0.2). Lower bias and toxicity scores ($\downarrow$) indicate better performance, while higher KR, Faith., and Rel. scores ($\uparrow$) suggest improved retention of useful information. Best scores are shown in bold. The original biased and debiased sentences (our safe ground truth label for debiasing) pre-safety interventions are derived from the original data. Post-Debiasing scores are the performance of LLMs that are fine-tuned or few-shot prompted.
@@
We evaluated the effectiveness of different LLMs in mitigating bias through language generation using this dataset portion, while preserving knowledge. Llama2-7B-chat and Mistral-7B-instruct-v0.2 are parameter efficient fine-tuned on this data (Hyperparameters in Table 20). Effectiveness is measured using metrics such as bias, toxicity, content retention, faithfulness, and answer relevancy, as defined by DeepEval [13] and detailed in Appendix G.6.3. Our review team verified the evaluations to ensure they meet expected standards.
Results: Table 8 shows that Mistral-7B (Fine-tuned) outperforms other models, achieving the lowest Bias and Toxicity and highest KR, Faith., and Rel. scores. Llama2-7B, both few-shot and fine-tuned, shows improvements but still compromises in some areas. Overall, fine-tuning safe LLMs like Llama2 or Mistral with our debiasing dataset further enhances their performance.
@@ 4 Conclusions
In this work, we present BEADs, a comprehensive dataset specifically labeled for bias detection across multiple NLP tasks, including text classification, token classification, bias quantification, and debiasing through benign language generation.The dataset is annotated using a two-phase approach, combining active learning on GPT-3.5 generated labels with GPT-4 providing higher-quality labels, which are then verified by expert annotators for reliability. Benchmarking the dataset using BERT-like models and autoregressive LLMs (Section 3), we find that smaller models, such as BERT and RoBERTa, outperform LLMs in bias classification tasks. Furthermore, we observe that some LLMs tend to perpetuate biases for certain demographics. However, fine-tuning LLMs on safe data from BEADs tends to reduce these biases while maintaining the models’ overall performance.
Limitations
While our study comprehensively addresses various aspects of bias, it is crucial to acknowledge the inherently subjective nature of bias itself. We implemented a two-phase approach using GPT-3.5 for preliminary labels and GPT-4 for higher-quality gold labels. This approach, while improving label quality, is a more complicated annotation process. Our evaluations utilized state-of-the-art baseline models and focused on straightforward prompt scenarios. Future research could explore the effects of adversarial or misleading prompts [78], which may further test and enhance the models’ detection capabilities. Despite careful review of our gold standard dataset, the potential for human errors and residual subjectivity remains.
In the empirical analysis between different models, we did not have access to the log-probability of the chat and instruct models. Consequently, the models were only evaluated on generative responses, which may be why these models showed lower performance. While we demonstrated performance across various NLP tasks, there remain areas for bias evaluation in NLP tasks, such as question answering, machine ethics, and translation that still need to be addressed.
Although we are confident in our evaluation due to reliable gold data labels, we recognize the need for further model fine-tuning for safety and benign language generation. Given that many institutions do not disclose detailed research information, such as model specifics, and evaluation scripts, we believe that incorporating these results for two main types of models (encoder-only and autoregressive LLMs) will provide a comparative view of the state-of-the-art moels and facilitate training more models of same family. We also recognize the risk of data misuse by malicious users who may exploit the unintended functionalities of language models. Therefore, it is crucial for model developers to take precautions and implement appropriate safeguards, such as carefully designed prompts and fine-tuning strategies, to ensure that these models are not used to cause harm For a more detailed discussion on limitations, refer to Section I.
Broader Impact
The insights from this research can inform the pre-training or safe training processes of LLMs to better handle and mitigate linguistic biases. Our work highlights vulnerabilities in dealing with diverse demographic data and can contribute to the development of responsible AI practices. Although our research may be limited in scope, it can inspire further questions in the growing area of LLM safety and AI risks. For a more detailed discussion, refer to Section J.
@@ Acknowledgements
Appendix A Literature Review
Recent advances in NLP, particularly with the deployment of LLMs like GPT-4 and ChatGPT, necessitate comprehensive benchmarks to address biases [75] and toxicity [16] in sentiment analysis [76, 44]. Studies on word embeddings have revealed gender biases [10], and creating balanced datasets has improved fairness scores across demographic groups [18]. Robust datasets such as Jigsaw [4], HolisticBias [69], and others [83, 70, 28, 7, 12, 22] support the mitigation of bias and toxicity, promoting comprehensive evaluation of models on fairness, safety, and robustness. Previous efforts like WinoGender [63], RedditBIAS [7], and BBQ [53], disability [69] and sexual orientation [25], across various NLP tasks. More details can be found in Appendix A.1. Building on the seminal work, we cover a number of NLP tasks on classification like bias, toxicity and sentiment, all in one dataset and we have explored a wide range of biases aspects [56] such as gender, race, and socioeconomic status.
Benchmarking on Bias Entity Extraction
Named Entity Recognition (NER), though not widely used method for bias identification, is crucial in identifying biases in NLP [57, 58]. Studies show NER models perform better for certain demographics [48], such as recognizing fewer female names as PERSON entities compared to male names in U.S. census data [46]. Some work uses NER to address biases in names and identities (persons, citites etc) by adding learnable adversarial noise to those identities [29]. Counterfactual data augmentation also mitigates gender bias in word embeddings [43, 80]. The NERD framework evaluates linked data entity extractors on semi-structured texts [59]. For some NER tasks, Bi-LSTM with a CRF layer, incorporating pre-trained embeddings like Glove and BERT, and a classification model for sentence splitting is used [51]. The entity debiasing framework for fake news detection models causal relationships among entities, contents, and veracity [88]. Despite various approaches, there is room for dedicated data preparation. We present a dataset identifying bias indicators for use in entity extraction tasks, allowing NLP models to detect biased named entities.
Benchmarking on Bias Quantification through Demographics
Benchmarking on Benign Language Generations
A.1 Existing Datasets Comparions with BEADs
Table 9: Comprehensive Comparison of Bias Datasets: This table presents a detailed comparison of various datasets used for bias analysis in NLP. Each entry provides the name of the dataset (Dataset), the original source of the data (Source), specific details including the size of the dataset, the style of annotation (Annot.), and accessibility information (Size/Annot./Open), the primary task for which the dataset is designed (Task), and the specific type of bias that the dataset aims to address or measure (Dimension). Acronyms used: O: Open access (✓for ‘yes’), H: human-labeled, quant.: quantification, eval.: evaluation, LM: language model, H-M: Human & Machine learning annotations.
@@ Appendix B Prompts for Annotation
B.1 Prompt for Initial Labeling
@@ B.1.3 Aspect of Bias Annotation Task
Appendix C Initial Data Labeling
@@
We used GPT-3.5 (gpt-3.5-turbo), equipped with few-shot prompts (2 examples), to generate initial labels for various features across over 30,000 samples. The labels generated are Biases, toxicity, sentiment, biased words, aspects, and demographic mentions.
@@
Outputs from BERT were reviewed, focusing on predictions with low confidence. Continuous improvement in data labeling accuracy and classifier performance through feedback and adjustment from our reviewers, details in Appendix E.
@@
Appendix D Gold Data Labeling

### v3

@@ 5.1 Text Classification for Bias, Toxicity, and Sentiment
We fine-tuned smaller BERT-like models for the task. We also fine-tuned autoregressive LLMs including Llama-2-7B-chat [72] and Mistral-7B-Instruct-v0.2 [3] using QLoRA [68], and evaluated them in standard and few-shot setups. Unlike BERT-like models that output probability scores, these LLMs produce categorical outputs, so we analyzed them for precision, recall, and accuracy as in [70].
Results: As shown in Table 5, fine-tuning smaller BERT-like models perform better than LLMs on specific bias classification tasks. This demonstrates their efficiency and practicality for real-world applications. Fine-tuned BERT-like models achieved approximately 85% accuracy in text classification, which is better than autoregressive LLMs in fine-tuning as well as few-shot setups.
The observed performance differences between smaller language models and LLMs ( LLMs have billions of parameters) could stem from the differing evaluation metrics used, such as probability scores versus confidence scores. Investigating these discrepancies might require a deeper exploration of factors like architectural variations, the sizes of training datasets, and the specific demands of the tasks they are designed to perform. The performance on out-of-distribution datasets, such as MBIB [73], Toxigen [21], and financial news [74] are reported in Table 21 and the results aligns with the findings discussed in this section.
@@ 5.2 Token Classification Task
For the token classification task, we fine-tuned smaller models including BiLSTM-CRF, BERT-like models, and LLMs like Llama2-7B-chat and Mistral2-7B-instruct. These models were evaluated on their ability to classify words as ‘Bias’ or ‘O’ (outside bias). We present this portion of data as CoNLL Bias Dataset following CONLL-2003 format [27] standards for NER, where ‘B-BIAS’ and ‘I-BIAS’ tags are consolidated into ‘Bias’, and ‘O’ into ‘Non-Bias’ for evaluation.
@@
Results: As shown in Table 6, the smaller fine-tuned BERT-like models achieved the best scores across all metrics, in classifying both ‘Non-Bias’ and ‘Bias’ entities. Few-shot models like Llama2-7B and Mistral-7B delivered moderate performance, where as fine-tuning LLMs tend to show increased performance for token classification task, compared to few-shots in prompt setting. In this experiment also, we observe that smaller language models perform better than LLMs in the token classification task. This difference in performances may be attributed to factors such as the efficiency of smaller models in processing and learning from limited data, their faster adaptation to specific tasks, or the potential overfitting issues in larger models with more parameters.
@@ 5.3 Bias Quantification on Demographics
Task Setting: This task evaluates the ability of a LLM to predict tokens that are demographic identifiers, such as gender, race, and religion in sentences. The intuition is that the model’s choices in replacing placeholders like GENDER, RACE, and RELIGION reveal potential biases. The test set includes both biased and neutral examples, and the task measures the rate at which the LLM associates certain demographics with biased or neutral content. Format: The sentences are formatted with placeholders for demographic indicators, structured as <text with identity mention>. For example, if we have a sentence template like “The {identity mention} manager is irritatingly bossy", labeled as Bias, and the LLM consistently fills it with “The female manager is irritatingly bossy" (tested five times, we applied majority here), this indicates a bias tendency towards such identity mentions. This test is performed using our Bias Quantification Demographics Dataset.
We evaluated both smaller BERT-like models and autoregressive LLMs by testing their ability to fill in placeholders. Their performance was assessed based on Bias Rate metric, we defined earlier. The comparison with ground truth bias labels is presented in Table 7 .
@@
Table 7: Bias Rate for Identity Mentions Across Different Models. Lower bias rate ($\downarrow$) is desirable. Models exacerbating highest bias towards a demographic are highlighted in red.
Identity Mention Original DistilBERT BERT RoBERTa Mistral-7B-v0.2 Ethnicity Black 66.7% 46.4% 53.8% 43.2% 50.0% Ethnicity White 57.1% 77.8% 57.1% 54.5% 50.0% Religion Christian 27.6% 33.3% 37.7% 51.1% 26.3% Religion Jewish 50.0% 63.3% 63.2% 59.6% 50.0% Religion Muslim 40.0% 56.6% 57.8% 55.7% 50.0% Gender Female 57.7% 59.7% 49.2% 57.1% 56.4% Gender Male 46.0% 52.7% 55.3% 53.3% 43.0%
@@
Task Setting: This task assesses the ability of LLMs to evaluate neutrality in sentences that reference various demographic groups. We have selected 520 neutral sentences from our dataset and modified them to include variations pertaining to gender (male, female), race (Asian, Black, Hispanic, White), and religion (Christian, Muslim, Hindu, Buddhist). The objective is to prompt the LLM to show whether these sentence variations exhibit bias or maintain neutrality with respect to the mentioned demographic group. Format: Entries are formatted by integrating demographic variations into a standard prompt, such as <prompt><text><demographic group variation>. For example, comparing the sentences “Asians are good at math….." versus “Europeans are good at maths…..", the task challenges the LLM to identify which statement, if any, is neutral or is biased. The dataset examples are in our Demographic Variation Dataset.
LLMs like Mistral-7B-Instruct-v0.2 and Llama2-7B-chat are used during inference to identify and respond to sentences with demographic variation in the prompts. The models responses to these prompts, which reveal their tendencies to misclassify neutral statements as biased or vice versa.
@@
Results: The analysis presented in Figure 3 indicates that both Llama2-7B-chat and Mistral-7B-Instruct-v0.2 generally classify neutral sentences as “Neutral” across most demographic categories, which suggests high accuracy. However, there are notable exceptions in categories like “Black”, “Hispanic", “Muslim", and “Christian", where a higher proportion of “Biased” responses might reflect the models heightened sensitivity to or inherent biases in these demographics. Discussion on the result of Stereotype Alignment task on Stereotype Prompts Dataset is in F.2.
@@ 5.4 Language Generation Task
Task Setting: This task generates neutral or benign versions of biased texts, the goal is to remove bias in language while retaining knowledge or context of the original sentence. The dataset portion has 8,300 entries with various demographics and bias aspects. Format: Each sentence pair is presented as <original sentence>, <benign variation>. For example, Original Sentence: “She lacks the aggressive approach needed for effective leadership in this role", Benign Sentence: “The candidate’s leadership style does not align with the assertive approach required for this role". The dataset of our dataset for this purpose is Language Generation Dataset. We also format this dataset portion into Alpaca [28] format for instruction-fine-tuning.
Table 8: Comparison of Bias, Toxicity, Knowledge Retention (KR), Faithfulness (Faith.), and Answer Relevancy (Rel.) across different model variations (Llama2-7B-chat and Mistral-7B-instruct-v0.2). Lower bias and toxicity scores ($\downarrow$) indicate better performance, while higher KR, Faith., and Rel. scores ($\uparrow$) suggest improved retention of useful information. The original biased and debiased sentences (our ground truth label for debiasing) pre-debiasing are derived from the original data. Post-Debiasing scores are the performance of LLMs that are fine-tuned or used with few-shot prompts.
@@
We evaluated the effectiveness of different LLMs in mitigating bias through language generation using this dataset portion. Llama2-7B-chat and Mistral-7B-instruct-v0.2 are used in few-shot settings, and also fine-tuned. Effectiveness is measured using LLM based metrics such as bias, toxicity, content retention, faithfulness, and answer relevancy [71] and our review team verified the samples of the evaluations to ensure they meet expected standards.
Results: Table 8 shows that fine-tuning LLMs in semi-supervised manner performs better in debiasing (shown with lower bias and toxicity scores) and maintain knowledge (with higher knowledge retention, faithfulness and relevancy scores). The fine-tuning results are better than few-shot prompting. In this experiment, zero-shot results are excluded, as preliminary trials demonstrated comparatively lower performance in those settings. The findings also indicate that LLMs fine-tuned on benign texts can effectively serve as debiasing agents, so we can have their better overall utility.
6 Discussion
Broader Impact
Our research into bias detection within datasets has many social implications, especially for the development and deployment of LLM applications. Our research pinpoints various biases present within datasets, which could potentially influence the training of language models. We provide methods for quantifying these biases and techniques for debiasing texts. The long term goal of this study is to develop applications that specifically support disadvantaged or marginalized groups, with a focus on promoting fairness and inclusivity in linguistic representation and beyond. We encourage the research community to utilize and expand upon our dataset by following the guidelines provided in our data construction framework.
Limitations
Subjective nature of bias and experimental nature of the labeling methods While our study covers multiple aspects of bias, we recognize that bias is inherently subjective and can vary greatly depending on context.
In this work, we used a two-phase approach: initially, we employed GPT-3.5 for preliminary labeling and then utilized GPT-4 for creating higher-quality, gold-standard labels. This choice was driven by the experimental nature of our work and the availability of newer model versions, which encouraged us to explore the latest advancements. However, we also acknowledge that the associated costs of using these models can be quite high and may not always be feasible for everyone.
Need for more prompting methods Our evaluations utilized state-of-the-art baseline models and focused on straightforward prompt scenarios. Future research could explore the effects of adversarial or misleading prompts [75], which may further test and enhance the models’ detection capabilities. Despite careful review of our gold standard dataset, the potential for human errors and residual subjectivity remains, as mentioned by contemporary research also [76, 40].
Evaluation of LLMs vs smaller language models In the empirical analysis between different models, we did not have access to the log-probability of the chat and instruct models. Consequently, these LLMs were only evaluated on generative responses, which may be why these models showed lower performance, compared to encoder only smaller language models.
Lack of available methods Given that many institutions do not disclose detailed research information, such as model specifics, and evaluation scripts, we believe that incorporating these results for two main types of models (encoder-only and autoregressive LLMs) will provide a comparative view of the state-of-the-art moels and facilitate training more models of same family. While we demonstrated performance across various NLP tasks, there remain areas for bias evaluation in NLP tasks, such as question answering, machine ethics, and translation that still need to be addressed.
Dataset usage risk We also recognize the risk of data misuse by malicious users who may exploit the unintended functionalities of language models. Therefore, it is important for model developers to take precautions and implement appropriate safeguards, such as carefully designed prompts and fine-tuning strategies, to ensure that these models are not used to cause harm.
@@ 7 Conclusions
In this work, we present BEADs, a comprehensive dataset specifically labeled for bias detection across multiple NLP tasks, including text classification, token classification, bias quantification, and debiasing through benign language generation. The dataset is annotated using a two-phase approach that combines active learning with GPT-3.5-generated labels and higher-quality labels provided by GPT-4, which are then verified by expert annotators for reliability. Benchmarking the dataset with BERT-like models and autoregressive LLMs, we find that smaller models, such as BERT and RoBERTa, outperform LLMs in bias classification tasks. Furthermore, we observe that some LLMs tend to perpetuate biases for certain demographics, with smaller language models being more prone to biases. However, fine-tuning LLMs on safe data from BEADs tends to reduce these biases while maintaining the models’ overall performance.
@@ Acknowledgements
CRediT authorship contribution statement
Shaina Raza: Conceptualization, Investigation, Formal analysis, Methodology, Project administration, Software, Experimentation, Validation, Visualization, Writing – original draft, Writing – review, & editing, Supervision. Mizanur Rahman: Investigation. Experimentation. Formal Analysis. Validation. Review. Michael R. Zhang : Validation, Writing – review& editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data and code is made available.
@@
Appendix A Comparison with Existing Datasets
Table 9: Comparison of Bias Datasets: Each entry includes the dataset name (Dataset), source (Source), dataset details, annotation style (Annot.: Human (H) or Model-based (M)), accessibility (Size/Annot./Open), primary task (Task), and bias type (Dimension). Acronyms: O: Open access (✓for ‘yes’), H: human-labeled, quant.: quantification, eval.: evaluation, LM: language model, H-M: Human & Model-based annotations.
@@ Appendix B Prompts for Annotation
B.1 Prompt for Initial Data Labeling
@@ B.1.3 Aspect of Bias Annotation Task
B.2 Initial Data Labeling and Active Learning
@@
We used GPT-3.5 (gpt-3.5-turbo), equipped with few-shot prompts (2 examples), to generate initial labels for various features across over 50,000 samples. The labels generated are Biases, toxicity, sentiment, biased words, aspects, and demographic mentions.
@@
Outputs from BERT were reviewed, focusing on predictions with low confidence. Continuous improvement in data labeling accuracy and classifier performance through feedback and adjustment from our reviewers, details in Appendix C.
@@
B.3 Gold Data Labeling